# Localized Band Gap Resonances in Defect-embedded Monoatomic Chains

### v1

@@ II.1 Defect eigenfrequency and eigenmode shape
represent the matrix entries of $\mathbf{K}_{\mathrm{d}}$. Eq. (3) can be solved numerically, assuming a large value of $n$ (i.e, large $\mathrm{N}$) to estimate $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Fig. 2(c-ii,iv) plot the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ contours, respectively, obtained using the analytical approach, showing an excellent correlation with their numerical counterparts in Fig. 2(c-i,iii). Note that the decay rate, $\kappa_{I}$ only depends on the base MoA properties $k_{\mathrm{g,r}}$ and defect frequency, $\Omega_{\mathrm{d}}$. Therefore, each $\Omega_{\mathrm{d}}$ is associated with a unique $\kappa_{I}$ leading to identical shapes for iso-defect frequency and iso-decay rate curves in Fig. 2c.
Alternately, for the bulk DMoA (Fig.1c), we utilize the U-transformation method[44] proposed by Cai et al.[32] to arrive at an implicit equation identical to Eq. (3) that can be solved numerically to estimate the defect frequency and mode shape. Though an explicit analytical form of $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ remains elusive, the proposed solution approach for both DMoAs require a numerical calculation only at the terminal step to solve a well-defined implicit analytical equation (Eq. (3)). Therefore, the analytical approach executed with a numerical algorithm[45, 46] of our choice allows a systematic derivation of DMoA material parameters that produce the desired spatio-temporal defect characteristics.
II.2 Velocity amplitude envelope at defect mode resonance
@@
A closer look at the iso-defect frequency and iso-decay rate curves in Fig. 2c indicates the presence of multiple $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ pairs that can manifest the same $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Therefore, we formulate and study an additional quantity associated with resonance dynamics at the defect location, i.e., the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ (see Sec. A3) to distinguish the dynamic behavior of edge DMoAs with distinct material properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. Absolute velocity (displacement) amplitudes at resonance have rarely been featured in fundamental research studies of AMMs. However, their emerging relevance in determining AMM efficiency for applications, e.g., energy harvesting[36, 37], flow control via fluid-metamaterial interactions[38, 42] motivates their consideration as a critical behavioral parameter describing the AMM dynamics. Hence, a prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ is considered in this section as the third and final parameter to complete the edge DMoA design.
@@
As anticipated, the mode shapes associated with DMoAs described by the $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers, plotted in Fig. 3b manifest an identical decay rate, $\kappa_{I}=0.962$. Each DMoA is now subject to a sinusoidal external force, $\mathrm{F}(\bar{t})=\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t})$ at the edge defect mass, and Fig. 3(c-i) plots the corresponding numerical and analytical velocity response,
@@
Consistent with a typical response at resonance, all DMoAs exhibit an oscillatory velocity response at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ with a linearly increasing velocity amplitude in time. However, each DMoA manifests a unique rate of increase in velocity amplitude that we designate as the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$. $\dot{\mathrm{V}}_{\mathrm{E}}$ can be analytically estimated in terms of the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ or $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, and the input force amplitude, $\mathrm{F}_{0}$ (see Sec.A3). Fig. 3(c-ii) contrasts the analytical resonance solution in Eq. (6) and the velocity response recorded in the numerical simulations ($\square$ - marker, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}=\{1,0.382\}$), showing an excellent correlation between the two. Therefore, the proposed analytical approach offers a broad set of design parameters to construct a customized DMoA based on operational requirements.
II.3 Multi-defect embedded grounded monoatomic chain
Figure 4: Multi-defect embedded MoA (Multi-DMoA). (a) Absolute Percentage error contours for a Multi-DMoA ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. (b) Edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, of a Multi-DMoA with an edge defect at $x_{\mathrm{d1}}=0$ and a bulk defect at $x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$. FRFs of a Multi-DMoA denoted by
@@ II.3 Multi-defect embedded grounded monoatomic chain
-marker in (a,b) when the (c-i) edge and (c-ii) bulk defect masses are excited and, their displacement response monitored, (c-iii) eigenmode shapes of the Multi-DMoA.
@@
In Sec. II.1 and Sec. II.2 we have established the accuracy of the proposed analytical approach in determining the dynamic characteristics of a DMoA with a single defect. But, what are the limitations of the analytical approach? Can the approach be used to characterize DMoAs with multiple defects? To address these questions and investigate the robustness of the solution procedure, in this section, we contrast the numerical and analytical dynamic characteristics of a multi-defect embedded grounded monoatomic chain (Multi-DMoA).
@@
Fig. 4a plots the absolute percentage error between the analytical and numerical values of the two defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}(>\Omega_{\mathrm{d1}})$ generated due to the presence of two defects at locations, $x_{\mathrm{d1}}$ and $x_{\mathrm{d2}}$ in a Multi-DMoA ($\mathrm{N}=21$). The diagonal direction in both these plots remains unpopulated as we assume distinct locations for the embedded material defects, i.e., $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. The sea of dark blue, low error values ($0.0005\%-2\%$) in both plots indicate the validity of the analytical approach derived in Sec. II.1 for a large section of the Multi-DMoA configurations provided, the two defects in the Multi-DMoA remain relatively isolated from each other. Alternately, as the defect locations become proximal, i.e., within 3 masses, marked by the off-diagonal lines, $x_{\mathrm{d2}}=x_{\mathrm{d1}}\pm 3$, the Multi-DMoA configuration deviates from the isolated (single) defect assumption giving rise to a region of high error ($>2\%$, max. error $\sim 33\%$). However, note that the expanse of the high error region is dependent on the decay rate, $\kappa_{I}$ associated with bulk defect mode, i.e., as the $\kappa_{I}$ grows larger, the high error region shrinks (Sec. A4). Therefore, a modified analytical approach is required to investigate the design space of proximal defects. This lies outside the scope of the current study and isn’t pursued further. Although, prior efforts have studied the dispersion characteristics of phononic crystals with multiple proximal defects via super-cell analysis[26], that can provide good qualitative insights into a potential solution approach.
@@
Another important observation to highlight in the $\Omega_{\mathrm{d1}}$ error plots in Fig. 4a, is the linear light blue regions - $x_{\mathrm{d1}}=1,\mathrm{N}-2$ and $x_{\mathrm{d2}}=1,\mathrm{N}-2$. These lines correspond to Multi-DMoA configurations where one of the penultimate masses is a defect. Technically this defect can be categorized as a bulk defect, however, the proximity to the finite boundary of the Multi-DMoA has a significant impact on the eigenmode characteristics, leading to a slightly higher error ($\sim 3.9\%$) compared to the Multi-DMoA configurations in the dark blue regions. The impact of boundary effects is less prevalent in cases where the defect frequencies lie in the $2^{\mathrm{nd}}$ BG of the MoA (see Sec. A4). We can attribute this phenomenon to the real part of the wavenumber, $\kappa_{R}$ of these defect modes. The defect mode in the $1^{\mathrm{st}}$ BG has $\kappa_{R}=0$ that requires the boundary adjacent to the defect location to vibrate in-phase with the defect at resonance, leading to a rigid body-like motion of the boundary and defect masses. Alternately, the $2^{\mathrm{nd}}$ BG requires $\kappa_{R}=\pi$, thus, allowing the boundary mass to vibrate out of phase with the defect leading to relatively decoupled, anti-symmetric dynamics between the sites rendering the actual defect frequency for the latter case closer to the analytical estimate than the former.

### v2

@@ 2.1 Defect eigenfrequency and eigenmode shape
represent the matrix entries of $\mathbf{K}_{\mathrm{d}}$. Eq. (3) can be solved numerically, assuming a large value of $n$ (i.e, large $\mathrm{N}$) to estimate $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Fig. 2(c-ii,iv) plot the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ contours, respectively, obtained using the analytical approach, showing an excellent correlation with their numerical counterparts in Fig. 2(c-i,iii). Note that the decay rate, $\kappa_{I}$ only depends on the base MoA PnC properties $k_{\mathrm{g,r}}$ and defect frequency, $\Omega_{\mathrm{d}}$. Therefore, each $\Omega_{\mathrm{d}}$ is associated with a unique $\kappa_{I}$ leading to identical shapes for iso-defect frequency and iso-decay rate curves in Fig. 2c.
Alternately, for the bulk DMoA PnC (Fig.1c), we utilize the U-transformation method[44] proposed by Cai et al.[32] to arrive at an implicit equation identical to Eq. (3) that can be solved numerically to estimate the defect frequency and mode shape. Though an explicit analytical form of $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ remains elusive, the proposed solution approach for both DMoA PnCs require a numerical calculation only at the terminal step to solve a well-defined implicit analytical equation (Eq. (3)). Therefore, the analytical approach executed with a numerical algorithm[45, 46] of our choice allows a systematic derivation of DMoA PnC material parameters that produce the desired spatio-temporal defect characteristics.
2.2 Velocity amplitude envelope at defect mode resonance
@@
A closer look at the iso-defect frequency and iso-decay rate curves in Fig. 2c indicates the presence of multiple $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ pairs that can manifest the same $\{\Omega_{\mathrm{d}},\kappa_{I}\}$. Therefore, we formulate and study an additional quantity associated with resonance dynamics at the defect location, i.e., the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ (see Sec. A3) to distinguish the dynamic behavior of edge DMoA PnCs with distinct material properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. Absolute velocity (displacement) amplitudes at resonance have rarely been featured in fundamental research studies of PnCs. However, their emerging relevance in determining PnC efficiency for applications, e.g., energy harvesting[36, 37], flow control[38, 42] motivates their consideration as a critical behavioral parameter describing the PnC dynamics. Hence, a prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ is considered in this section as the third and final parameter to complete the edge DMoA PnC design.
@@
As anticipated, the mode shapes associated with DMoA PnCs described by the $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers, plotted in Fig. 3b manifest an identical decay rate, $\kappa_{I}=0.962$. Each DMoA PnC is now subject to a sinusoidal external force, $\mathrm{F}(\bar{t})=\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t})$ at the edge defect mass, and Fig. 3(c-i) plots the corresponding numerical and analytical velocity response,
@@
Consistent with a typical response at resonance, all DMoA PnCs exhibit an oscillatory velocity response at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ with a linearly increasing velocity amplitude in time. However, each DMoA PnC manifests a unique rate of increase in velocity amplitude that we designate as the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$. $\dot{\mathrm{V}}_{\mathrm{E}}$ can be analytically estimated in terms of the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$ or $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, and the input force amplitude, $\mathrm{F}_{0}$ (see Sec.A3). Fig. 3(c-ii) contrasts the analytical resonance solution in Eq. (4) and the velocity response recorded in the numerical simulations ($\square$ - marker, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}=\{1,0.382\}$), showing an excellent correlation between the two. Therefore, the proposed analytical approach offers a broad set of design parameters to construct a customized DMoA PnC based on operational requirements.
2.3 Damped DMoA PnCs
The analytical approach described above can be extended to DMoA PnCs containing dissipative elements, by taking into account the damping matrix, $\mathbf{C}$ and modifying the dynamic stiffness matrix, $\mathbf{K}_{\mathrm{d}}=\mathbf{K}+i\Omega\mathbf{C}-\Omega^{2}\mathbf{M}$, resulting in matrix coefficients:
$d=2+k_{\mathrm{g,r}}+i\Omega(c_{\mathrm{on,r}}+2c_{\mathrm{in,r}})-\Omega^{2},a=-(1+i\Omega c_{\mathrm{in,r}}),\beta=-(1+i\Omega c_{\mathrm{in,r}}),$
$\alpha=(k_{\mathrm{gd,r}}-k_{\mathrm{g,r}})-(1+i\Omega c_{\mathrm{in,r}})+(1-m_{\mathrm{r}})\Omega^{2};$
where, the general damping matrix, $\mathbf{C}=c_{\mathrm{on,r}}\mathbf{I}+c_{\mathrm{in,r}}\mathbf{K}_{\mathrm{in}}$ (see Sec. A1), accommodates both onsite and intersite damping elements - $c_{\mathrm{on,r}}=c_{\mathrm{on}}/\sqrt{km}$ and $c_{\mathrm{in,r}}=c_{\mathrm{in}}/\sqrt{km}$ , respectively. Substituting the above coefficients into, Eq. (3) will provide the defect mode frequencies and mode shapes in the damped DMoA PnCs.
In addition, in the presence of damping ($c_{\mathrm{on,r}}\neq 0$ and/or $c_{\mathrm{in,r}}\neq 0$), the DMoA PnC produces a bounded steady-state response,
$\dot{v}_{\mathrm{d}}=\dot{\mathrm{V}}_{\mathrm{SS}}\left[\alpha\Omega_{\mathrm{d}}\sin(\Omega_{\mathrm{d}}\bar{t})+\cos(\Omega_{\mathrm{d}}\bar{t})\right],$
at the edge defect, when subject to external forcing, $\mathrm{F}(\bar{t})=\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t})$. Therefore, the steady-state velocity amplitude, $\dot{\mathrm{V}}_{\mathrm{SS}}=-b_{I}\Omega_{\mathrm{d}}\widetilde{\mathrm{V}}_{\mathrm{d}}$ (Sec. A4) can be utilized as an alternate defect mode characteristic to the resonance velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ in damped DMoA PnCs.
2.4 Multi-defect embedded grounded monoatomic phononic crystal
Figure 4: Multi-DMoA PnCs. (a) Absolute Percentage error contours for a Multi-DMoA PnC ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. (b) Edge and bulk defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}$, respectively, of a Multi-DMoA PnC with an edge defect at $x_{\mathrm{d1}}=0$ and a bulk defect at $x_{\mathrm{d2}}\in[1,\mathrm{N}-2]$. FRFs of a Multi-DMoA PnC denoted by
@@ 2.4 Multi-defect embedded grounded monoatomic phononic crystal
-marker in (a,b) when the (c-i) edge and (c-ii) bulk defect masses are excited and, their displacement response monitored, (c-iii) eigenmode shapes of the Multi-DMoA PnC.
@@
In Sec. 2.1 and Sec. 2.2 we have established the accuracy of the proposed analytical approach in determining the dynamic characteristics of a DMoA PnC with a single defect. But, what are the limitations of the analytical approach? Can the approach be used to characterize DMoA PnCs with multiple defects? To address these questions and investigate the robustness of the solution procedure, in this section, we contrast the numerical and analytical dynamic characteristics of a multi-defect embedded grounded monoatomic phononic crystal (Multi-DMoA PnC).
@@
Fig. 4a plots the absolute percentage error between the analytical and numerical values of the two defect frequencies, $\Omega_{\mathrm{d1}}$ and $\Omega_{\mathrm{d2}}(>\Omega_{\mathrm{d1}})$ generated due to the presence of two defects at locations, $x_{\mathrm{d1}}$ and $x_{\mathrm{d2}}$ in a Multi-DMoA PnC ($\mathrm{N}=21$). The diagonal direction in both these plots remains unpopulated as we assume distinct locations for the embedded material defects, i.e., $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$. The sea of dark blue, low error values ($0.0005\%-2\%$) in both plots indicate the validity of the analytical approach derived in Sec. 2.1 for a large section of the Multi-DMoA PnC configurations provided, the two defects in the Multi-DMoA PnC remain relatively isolated from each other. Alternately, as the defect locations become proximal, i.e., within 3 masses, marked by the off-diagonal lines, $x_{\mathrm{d2}}=x_{\mathrm{d1}}\pm 3$, the Multi-DMoA PnC configuration deviates from the isolated (single) defect assumption giving rise to a region of high error ($>2\%$, max. error $\sim 33\%$). However, note that the expanse of the high error region is dependent on the decay rate, $\kappa_{I}$ associated with bulk defect mode, i.e., as the $\kappa_{I}$ grows larger, the high error region shrinks (Sec. A5). Therefore, an alternate approach is required to investigate the design space of proximal defects. This lies outside the scope of the current study. Although, prior efforts have studied PnCs with proximal defects via super-cell analysis[26, 47, 48, 49], that can provide good qualitative insights into a potential solution approach. Since we are primarily interested in localized BG resonances, analyzing an appropriately sized super-cell around the proximal defects, will reveal defect modes whose frequency and shape correlate to the complete PnC counterpart, as masses far away from the defect location remain (virtually) undeformed at defect resonance.
@@
Another important observation to highlight in the $\Omega_{\mathrm{d1}}$ error plots in Fig. 4a, is the linear light blue regions - $x_{\mathrm{d1}}=1,\mathrm{N}-2$ and $x_{\mathrm{d2}}=1,\mathrm{N}-2$. These lines correspond to Multi-DMoA PnC configurations where one of the penultimate masses is a defect. Technically this defect can be categorized as a bulk defect, however, the proximity to the finite boundary of the Multi-DMoA PnC has a significant impact on the eigenmode characteristics, leading to a slightly higher error ($\sim 3.9\%$) compared to the Multi-DMoA PnC configurations in the dark blue regions. The impact of boundary effects is less prevalent in cases where the defect frequencies lie in the $2^{\mathrm{nd}}$ BG of the MoA PnC (see Sec. A5). We can attribute this phenomenon to the real part of the wavenumber, $\kappa_{R}$ of these defect modes. The defect mode in the $1^{\mathrm{st}}$ BG has $\kappa_{R}=0$ that requires the boundary adjacent to the defect location to vibrate in-phase with the defect at resonance, leading to a rigid body-like motion of the boundary and defect masses. Alternately, the $2^{\mathrm{nd}}$ BG requires $\kappa_{R}=\pi$, thus, allowing the boundary mass to vibrate out of phase with the defect leading to relatively decoupled, anti-symmetric dynamics between the sites rendering the actual defect frequency for the latter case closer to the analytical estimate than the former.
Figure 5: 2D Multi-DMoA PnC. (a-i) 2D MoA PnC square lattice, the primitive unit cell, and an equivalent axis-symmetric configuration of a mass at $(x_{p},x_{q})$. (a-ii) A finite 2D DMoA PnC square lattice ($\mathrm{N}_{p}=\mathrm{N}_{q}=21$) with a corner (magenta), edge (blue), and a bulk (green) defect. (b-i) 2D dispersion curve (Eq. (6)) traversing the irreducible Brillouin zone $\Gamma-\mathrm{X}-\mathrm{M}-\Gamma$, i.e., $\kappa_{p}=0\rightarrow\pi,\kappa_{q}=0;\ \kappa_{q}=0\rightarrow\pi,\kappa_{p}=\pi;\ \kappa_{p}=\kappa_{q}=\pi\rightarrow 0$. (b-ii) Defect frequency and (b-iii) decay rate predicted by Eq. (10b) contrasted with the BG eigenmodes of the finite lattice in (a-ii). (c) Localized eigenmodes of the corner, edge, and bulk defect resonances. ($m_{\mathrm{r}}=4,k_{\mathrm{r}}=1,k_{\mathrm{g,r}}=4,k_{\mathrm{gd,r}}=0$)